OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has a length of square root of 128 feet and a width of square root of 200 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

1. \[\text{Perimeter} = 36 \sqrt{2} \text{ feet}\]2. The semi-perimeter is length plus width.\[\sqrt{128}+\sqrt{200} \]The perimeter is twice the semi-perimeter.\[\text{Perimiter}=2\left(\sqrt{128}+\sqrt{200}\right)=2\left(8 \sqrt{2}+10 \sqrt{2}\right)=2\left(18\sqrt{2}\right)=36 \sqrt{2} \]3. A floor.